# Premerger phenomena in neutron-star binary coalescences

### v1

@@ IV.1 Tides: general theory
Within the literature, one can find a few different definitions for the qualifiers “adiabatic”, equilibrium”, and “dynamical” in the context of tides. For example, following Yu et al. 436, the word “adiabatic” emphasises that one considers deformations in the zero-frequency limit $\Omega_{\rm orb}\rightarrow 0$. This is more typically called the equilibrium tide, though Yu et al. 436 make the distinction by allowing the latter to include $\Omega_{\rm orb}>0$ influences. Finite-frequency terms involve both non-resonant and resonant excitations, the latter of which these authors reserve for “dynamical” tide status. In other works, the inclusion of non-resonant, but time-dependent terms leads to corrections of the equilibrium quality factors which are then called “effective” [e.g. 308]. In this Review, the word “dynamical” conveys $\Omega_{\rm orb}>0$, though we are mostly interested in resonances and their connections to high-energy phenomena. We emphasise at this stage that many of the formulae that we present are done so in a Newtonian language. When one reaches the level of wishing to actually compare results with data in a serious way, it will generally be necessary to consider fully GR expressions insofar as possible [^4] Subtle issues arise in GR, making this non-trivial. For example, even the notion of eccentricity is ambiguous [208]. An all-encompassing notion of “tide” is also hard to define via the Weyl tensor or other means; see Section I.C in Pitre and Poisson 307.. Given that our main purpose here is pedagogical, we avoid writing out formalisms in this way, though point the interested reader toward relevant literature where appropriate. A depiction of these two class of tides is given in Figure 6.
@@
Following Lai and Wu 232 and others, we orient a spherical coordinate system around (the centre of mass of) $M_{A}$ with the “$z$-axis” directed along L. The gravitational potential produced by $M_{B}$ can then be expanded in terms of spherical harmonics $Y_{\ell m}$ [e.g. 315, 327]
@@
where $\mathcal{D}^{(\ell)}_{mm^{\prime}}$ is the so-called Wigner $\mathcal{D}$-function [e.g. 187, 429, 224], and $\phi_{s}=\phi+|\boldsymbol{\Omega}_{s}|t$.
@@ IV.1.1 Mathematical description and calculation methods
Modes with $n>1$ are referred to as “overtones”, as $n$ is defined by counting the number of radial nodes in the eigenfunction [see, e.g. 404]. Rotation, however, generally makes a separation of variables as above impossible, and a further sum over a dummy azimuthal index becomes necessary unless one further sets up a hierarchy in powers of $|\boldsymbol{\Omega}_{s}|$ [e.g. 365].
@@
In solving mode problems one must also take care to note that complex conjugates generally also solve the master equation, though we ignore this complication here for pedagogical purposes. [see 308, for a recent discussion].
At this stage, the amplitude of the modes have not yet entered, as these fall out of the homogeneous equation (17). Enter the tides. Formally this amounts to instead solving the inhomogenous version of the master equation,
@@
Fortunately, except possibly at very late stages in the inspiral, the tidal distortion of the spectrum is weak [113], as can be formally estimated with the Unno et al. 404 formula; see Sec. IV.2. One thus considers a “volume-averaged” problem where the amplitude evolution, $q_{\alpha}(t)$, of each mode is considered in isolation. The problem is thus reduced from 1+3 to 1+0 dimensions, and we will end up with some ODEs for the amplitude evolution. The key step involves introducing an inner-product,
@@
where the overlap integrals [^6] Two ways of generalising overlaps to GR have been proposed in Kuan et al. 224 and Miao et al. 267. The orthogonality between modes and the sum rule for tidal overlaps [326] are equally respected by both definitions, while only the former predicts a vanishing $g$-mode overlap in the zero stratification limit (at least for a simple, constant $\delta$ law; see Sec. III.3). This issue is related to overlap “leakage” discussed in Sec. VI.4.1. Throughout we adopt the definition of Kuan et al. 224. read
@@
Table 2: A non-exhaustive list of Post-Newtonian effects that influence the GW inspiral for compact objects. Constraints on each of these orders can be found in the violin plots produced by the LVK consortium; see Abbott et al. 8. In principle, stellar structure corrections and additional couplings occur at ever-higher orders.
@@ 0 Energy deposited into modes Alexander 16, 17
0.5 Scalar contributions to dynamics (non-GR) Yagi et al. 431
1 Stellar structure; pericentre advance Wagoner and Malone 416, Andersson et al. 27
1.5 Spin-orbit coupling; tail backscatter Blanchet et al. 65
2 Self-spin, spin-spin, quadrupole-monopole couplings Blanchet et al. 65, Poisson 309
2.5 GW back-reaction Schäfer 340, Blanchet et al. 66
3 Gravitational tails of tails Blanchet 63
4 Dissipative tidal number Ripley et al. 331, Saketh et al. 338
5 Gravitoelectric quadrupole Love number Damour et al. 106, Yagi 430
$6$ Gravitomagnetic quadrupole Love number Damour and Nagar 104, Henry et al. 178
$6.5$ Spin-tidal coupling Abdelsalhin et al. 10, Castro et al. 77
Thus, provided the free mode spectrum can be constructed, one need “only” solve equation (26). This, however, is still not quite the end of the story as the orbital phase $\Phi$ must be evolved simultaneously. By modelling the fluid motion in the neutron star as a set of harmonic oscillators [16, 17, 341] and incorporating mode kinetics into the Hamiltonian of the binary, Kokkotas and Schafer 211, Kuan et al. 224 have shown how this can be achieved with a high-order PN method, though with the modes themselves calculated in full GR without Cowling (i.e. including the metric perturbations). While not all are taken into account [such as those occurring at negative orders in gravities with non-metric fields where dominant mono- or dipolar radiation can exist; 431], we list PN orders at which various, important effects occur in Table 2.
PN effects in shaping stellar structure have been examined in several references [e.g. 416, 369, 27] and starts already at first order, though no systematic analysis of how these directly influence the inspiral has been carried out. Tides and spins contribute hierarchically at several PN orders accounting, for instance, for the breakdown of the point-particle approximation. Tides absorb orbital energy in the Newtonian theory already (i.e. at $0^{\rm th}$ order), while spin couples to the orbital angular momentum at the 1.5th order and to self-spin and companion spin at 2nd order. At 2.5th order enters the leading-order dissipation due to GW emission (cf. equation 10). The electric- (magnetic-) type stationary deformations factor into the dynamics at 5th (6th) order, as introduced in Sec. IV.1.2, and the coupling between this deformation to the spin follows at the next half order. GW170817 placed constraints on each of these PN orders, as detailed in Abbott et al. 8.
With the above in mind, the notion of dephasing can be made precise. In the stationary phase approximation, the frequency-space GW phase, $\Psi$, can be written as [e.g. 141, 101]
@@ IV.1.2 Tides: equilibrium
The equilibrium tide simply corresponds to the “$\Phi(t)=0$” portion of the dynamics. In this case, the quasiharmonic time dependence of (12) drops out, and our interest shifts to bulk, geometric deformations of the stellar surface (“zero-frequency oscillations”). The extent to which the stellar interior is suspectible to a time-independent tidal potential can be encapsulated by the (shape) Love numbers; effectively, much like the $Q_{\alpha}$ defined previously, these measure the extent of orthogonality between the stellar fluid (or solid) and some angular portion of the tidal field (i.e. $\ell,m^{\prime}$). In general, therefore, there is an infinite set of Love numbers, though often in the literature one will find the term “Love number” just to mean the quadrupolar, $m^{\prime}=0$ Love number. In the context of compact binaries in GR, Damour 103 first quantified how tidal Love numbers influence inspiral. Note also that if the star is static and axisymmetric at background order, the index $m^{\prime}$ falls out of the equations and thus the static, or even effective, Love numbers are characterised by a single harmonic number $\ell$. These reduced Love numbers are typically denoted $k_{\ell}$.
@@
which is the same $U$ from (12). Note that the multipole moments $I_{\ell m}$ are also implicitly defined at some PN order; see Section V in Thorne 396. At Newtonian order, $I_{\ell m}$ is defined through an integral over the mass-density weighted by a spherical harmonic and radius to power $\ell$:
@@
where the (axisymmetric) mass ($M_{\ell}$) and current ($S_{\ell})$ multipole moments can be defined either via the Thorne 396 or Geroch-Hansen [176] formulae. These formulations are totally equivalent by a theorem due to Gürsel 172. In the above we have the odd and even parity sectors of contractions of the trace-free Weyl tensor [e.g. 107, 291]. Using the schemes introduced by Pappas and Sotiriou 293, Suvorov and Melatos 383 to generalise the Geroch-Hansen definitions to theories beyond GR, one could try to extend the correspondence (37) to some other theory of gravity. This has not yet been attempted.
Figure 7: The main radial gravitational potential $A(R)$ — the lapse function of the effective-one-body spacetime over which the dynamics of a test particle describes the dynamics of the real binary system [70, 71] — as a function of the radial coordinate $R=a$ with $M=M_{A}+M_{B}$ (for this Figure only) in a tidal-EOB model. The binary black hole sector of this model (solid black) shows a deviation from the lapse of the Schwarzschild spacetime (dotted), indicating the dynamics of a test particle on the effective spacetime is different from that of a (plunging) orbit around a Schwarzschild black hole due to the finite symmetric mass ratio $\nu=M_{A}M_{B}/M^{2}$ of the binary (which should not be confused with the our use of $\nu$ as stellar spin used elsewhere). For the neutron star case, the gravitational potential including up to next-to-next-to-leading-order tidal effects (i.e. up to 7 PN; green dashed) is shown together with the result resumming the gravitational-self-force information up to 7.5 PN [60, red]. The specific equal-mass neutron-star binary considered here has a tidal deformability (equation 38) of $\kappa_{2}^{T}\approx 73.55$. From Bernuzzi et al. 54 with permission.
Tides effectively enhance the “attraction” between the components of the binary, leading to earlier merger relative to binary black-holes or point-particles. The tidal dynamics are elegantly described by a radial interaction potential in the effective-one-body (EOB) formalism [61, 54, 184]. PN models of tides in the EOB formalism were initially verified against numerical simulations by Baiotti et al. 42, Bernuzzi et al. 55, who demonstrated that the models can be improved with such calibrations but that it is not possible to accurately describe tides close to merger with a PN model. Such a restriction was lifted by Bernuzzi et al. 54 through resumming techniques. A schematic depiction of this attractive feature is shown in Figure 7, where the red curve describes the radial potential for a neutron-star binary with tidal resumming. A significant drop in the potential occurs at larger radius, indicating the stronger attraction felt by the binary than the equivalent binary black hole. The dominant term of the tidal imprints on waveform is delivered by the quadrupole quantity [106]
@@
for $\Lambda_{A}=\frac{2}{3}k_{2}^{A}(R_{A}/M_{A})^{5}$ with the same definition for object $B$ [see also 182, 142, 183, 105]. Note $k_{2}$ is defined in expression (33), effectively being the constant of proportionality between the quadrupole moment and the quadrupolar tidal potential (appropriately generalised to GR via the Weyl tensor; though cf. Footnote 4).
@@ IV.1.3 Tides: dynamical
Pulsation modes inside compact stars, tidally-forced or otherwise, are generally characterised according to the nature of the restoring force that ultimately damps the oscillations. The theory of mode oscillations is an active and rich area of research, which we cannot do justice in this Review. We therefore refer the reader to, e.g. Andersson and Kokkotas 30, Kokkotas et al. 210, Pratten et al. 313, Andersson 23, though a few QNM groups are particularly important in the context of GWs and tides, so we introduce them briefly.
@@
Modes that are instead restored by buoyancy are referred to as $g$-modes (i.e. “gravity” modes, not to be confused with GWs). The $g$-modes, in contrast to $f$-modes, do depend sensitively on the internal composition of the star (see Sec. III.3) and thus may be able to reveal a different kind of information. Depending on the chemical composition of the star, the $g$-modes tend to follow different relations as described by Kuan et al. 222. That is to say, whether the EOS is purely hadronic or hybrid (for instance) has a strong impact on the scaling of the modes with the mean density, temperature, and other, microphysical parameters. Generally speaking, the $g$-mode spectrum will be influenced by both composition and temperature (entropy) gradients. The former can be calculated from a EOS that provides the speed of sound and $d\epsilon/dp$ in a tabulated form so that the Brunt-Väisälä frequency can be determined [11, 404].
@@
The $p$-modes, and $f$-mode in particular, are especially important for astrophysical processes involving neutron stars owing to their compactness, including the tides: it can be shown that the $f$-mode couples most strongly to the tidal potential $\nabla U$ out of all other modes (at least for astrophysical stars, e.g. without near break-up rotation rates or virial-strength magnetic fields). The $f$-modes tend to have a (linear) frequency in the range of $2\lesssim\omega_{f}/2\pi\lesssim 3$kHz depending on the EOS [see 212, for a review], and thus are primarily relevant in the late stages of an inspiral. The dephasing induced by the $f$-mode oscillation is a problem that has attracted considerable attention, where the effects are usually incorporated via the Love number as an effective dressing [184, 360, 247, 34, 361, 153]. Considering rapidly rotating stars leads to a more complicated picture, since the retrograde mode may actually come into resonance much earlier as the inertial-frame frequency of the mode is $\omega_{i}=\omega_{0}-m\Omega_{s}$ (equation 23; see Kuan and Kokkotas 220, 221).
@@
In Kuan et al. 224, the method presented in Kokkotas and Schafer 211 to simultaneously evolve the inspiral and mode amplitudes was extended to 3.5 PN order. In particular, the conservative dynamics, leading-order gravitational radiation, and mode excitations (see Tab. 2) are incorporated into a total system Hamiltonian. Solving the associated equations of motion allows one to quantify the above effects, including that of spin-modulations in mode excitation (deferred to Sec. IV.3). The aforementioned physics can also be incorporated into the effective Hamiltonian of the EOB formalism [184, 360], where spin can also be included [361]. We note that only modes whose pattern speed is along the orbital motion can be considerably excited while there is another set of modes rotating in the opposite direction. Hereafter, we always discuss modes belonging to the former class unless stated otherwise.
@@
Figure 8: Left panel: self-consistent mode amplitude evolutions of $f$- and $g_{1,2,3}$-modes due to tidal forcing, together with the dephasing (lower; see equation 30) associated with the $f$- (magenta) and $g$-modes (other colors; see plot legends) as functions of time relative to the merger, $t-t_{\rm mrg}$. The GW frequency ($f_{\rm gw}$) is overplotted as the green line, corresponding to the vertical axis on the right-hand axis. Right: Same information as per the left panel though instead as functions of $f_{\rm gw}$, with $t-t_{\rm mrg}$ overlaid as the green line, corresponding to the vertical, right-hand axis. The evolution is carried for a $1.35+1.35\,M_{\odot}$ binary neutron-star (i.e. $q=1$) with the generalised piecewise-polytropic approximation of the APR EOS from O’Boyle et al. 282 (the original version of the piecewise-polytropic approximation in Read et al. 325 has discontinuous $\tilde{\Gamma}$, preventing an accurate determination of $g$-modes’ tidal properties when $\delta$ is small) with a constant stratification $\delta=0.005$. The vertical lines mark the frequency of $g_{1,2,3}$-modes, which are 90, 60 and 48 Hz, respectively.
@@
On the other hand, if the neutron star has a spin, the frequencies of retrograde modes will be reduced [98, 237, 435, 152], pushing the onsets of resonance earlier in the inspiral. Of the most relevant effects in terms of waveform is the $f$-mode excitation in a neutron star possessing an anti-aligned spin with the orbit [187, 247, 361], as also demonstrated in numerical simulations [e.g. 127, 153]; effects of spin are described in more detail in Sec. IV.3.
Current analytic waveform models include either implicitly or explicitly dynamical tides to introduce some effective dressing. In particular, Andersson and Pnigouras 33 show that the Love number can be expressed as a sum of the contributions of various mode with the $f$-mode being the primary source. As the orbital frequency increases, the tidal interaction will enter a dynamical regime, where the Love number varies with tidal frequency and is often referred to as an effective Love number [184, 360]. Another stream to model the dynamical tidal effects is to introduce amplification factors to the leading-order tidal effects to blend in higher-order PN contributions [105, 61, 277, 13]. Phenomenological fittings to numerical simulations have also been developed [119, 122, 121, 4] [see also 424].

### v2

@@ IV.1 Tides: general theory
Within the literature, one can find a few different definitions for the qualifiers “adiabatic”, equilibrium”, and “dynamical” in the context of tides. For example, following Yu et al. 441, the word “adiabatic” emphasises that one considers deformations in the zero-frequency limit $\Omega_{\rm orb}\rightarrow 0$. This is more typically called the equilibrium tide, though Yu et al. 441 make the distinction by allowing the latter to include $\Omega_{\rm orb}>0$ influences. Finite-frequency terms involve both non-resonant and resonant excitations, the latter of which these authors reserve for “dynamical” tide status. In other works, the inclusion of non-resonant, but time-dependent terms leads to corrections of the equilibrium quality factors which are then called “effective” [e.g. 314]. In this Review, the word “dynamical” conveys $\Omega_{\rm orb}>0$, though we are mostly interested in resonances and their connections to high-energy phenomena. We emphasise at this stage that many of the formulae that we present are done so in a Newtonian language. When one reaches the level of wishing to actually compare results with data in a serious way, it will generally be necessary to consider fully GR expressions insofar as possible [^4] Subtle issues arise in GR, making this non-trivial. For example, even the notion of eccentricity is ambiguous [213]. An all-encompassing notion of “tide” is also hard to define via the Weyl tensor or other means; see Section I.C in Pitre and Poisson 313.. Given that our main purpose here is pedagogical, we avoid writing out formalisms in this way, though point the interested reader toward relevant literature where appropriate. A depiction of these two class of tides is given in Figure 6.
@@
Following Lai and Wu 238 and others, we orient a spherical coordinate system around (the centre of mass of) $M_{A}$ with the “$z$-axis” directed along L. The gravitational potential produced by $M_{B}$ can then be expanded in terms of spherical harmonics $Y_{\ell m}$ [e.g. 321, 333]
@@
where $\mathcal{D}^{(\ell)}_{mm^{\prime}}$ is the so-called Wigner $\mathcal{D}$-function [e.g. 190, 434, 230], and $\phi_{s}=\phi+|\boldsymbol{\Omega}_{s}|t$.
@@ IV.1.1 Mathematical description and calculation methods
Modes with $n>1$ are referred to as “overtones”, as $n$ is defined by counting the number of radial nodes in the eigenfunction [see, e.g. 409]. Rotation, however, generally makes a separation of variables as above impossible, and a further sum over a dummy azimuthal index becomes necessary unless one further sets up a hierarchy in powers of $|\boldsymbol{\Omega}_{s}|$ [e.g. 370].
@@
In solving mode problems one must also take care to note that complex conjugates generally also solve the master equation, though we ignore this complication here for pedagogical purposes. [see 314, for a recent discussion].
At this stage, the amplitude of the modes have not yet entered, as these fall out of the homogeneous equation (18). Enter the tides. Formally, their accounting amounts to instead solving the inhomogenous version of the master equation,
@@
Fortunately, except possibly at very late stages in the inspiral, the tidal distortion of the spectrum is weak [114], as can be formally estimated with the Unno et al. 409 formula; see Sec. IV.2. One thus considers a “volume-averaged” problem where the amplitude evolution, $q_{\alpha}(t)$, of each mode is considered in isolation. The problem is thus reduced from 1+3 to 1+0 dimensions, and we will end up with some ODEs for the amplitude evolution. The key step involves introducing an inner product,
@@
where the overlap integrals [^6] Two ways of generalising overlaps to GR have been proposed in Kuan et al. 230 and Miao et al. 273. The orthogonality between modes and the sum rule for tidal overlaps [332] are equally respected by both definitions, while only the former predicts a vanishing $g$-mode overlap in the zero stratification limit (at least for a simple, constant $\delta$ law; see Sec. III.3). This issue is related to overlap “leakage” discussed in Sec. VI.4.1. Throughout we adopt the definition of Kuan et al. 230. read
@@
Table 2: A non-exhaustive list of Post-Newtonian effects that influence the GW inspiral for compact objects. Constraints on each of these orders can be found in the violin plots produced by the LVK consortium; see Abbott et al. 8. In principle, stellar-structure and multipole moments corrections and additional couplings enter at ever-higher orders [401].
@@ 0 Energy deposited into modes Alexander 16, 17
0.5 Scalar contributions to dynamics (non-GR) Yagi et al. 436
1 Stellar structure; pericentre advance Wagoner and Malone 421, Andersson et al. 28
1.5 Spin-orbit coupling; tail backscatter Blanchet et al. 66
2 Self-spin, spin-spin, quadrupole-monopole couplings, mag. dipoles Poisson 315, Ioka and Taniguchi 197
3 Gravitational tails of tails Blanchet 64
4 Dissipative tidal number Ripley et al. 337, Saketh et al. 344
5 Gravitoelectric quadrupole Love number Damour et al. 107, Yagi 435
$6$ Gravitomagnetic quadrupole Love number Abdelsalhin et al. 10, Henry et al. 181
$6.5$ Spin-tidal coupling Abdelsalhin et al. 10, Castro et al. 78
Thus, provided the free mode spectrum can be constructed, one need “only” solve equation (27). This, however, is still not quite the end of the story as the orbital phase $\Phi$ must be evolved simultaneously. By modelling the fluid motion in the neutron star as a set of harmonic oscillators [16, 17, 347] and incorporating mode kinetics into the Hamiltonian of the binary, Schäfer 346, Kokkotas and Schafer 217, Kuan et al. 230 have shown how this can be achieved with a high-order PN method, though with the modes themselves calculated in full GR without Cowling (i.e. including the metric perturbations). While not all are taken into account [such as those occurring at negative orders in gravities with non-metric fields where dominant mono- or dipolar radiation exists; 436], we list PN orders at which various effects occur in Table 2. We remark in this respect that although, strictly speaking, radiation-reaction is a 2.5 PN effect, often this is said to appear at Newtonian order when using a Newtonian definition for the system’s quadrupole moment(s) (see below). There is some ambiguity therefore in what constitutes a “PN order” exactly.
PN effects in shaping stellar structure have been examined in several references [e.g. 421, 374, 28] and starts already at first order, though no systematic analysis of how these directly influence the inspiral has been carried out. Tides and spins contribute hierarchically at several PN orders accounting, for instance, for the breakdown of the point-particle approximation. Tides absorb orbital energy in the Newtonian theory already (i.e. at $0^{\rm th}$ order), while spin couples to the orbital angular momentum at the 1.5th order and to self-spin and companion spin at 2nd order. At 2.5th order enters the leading-order dissipation due to GW emission (cf. equation 11). The electric- (magnetic-) type stationary deformations factor into the dynamics at 5th (6th) order, as introduced in Sec. IV.1.2, and the coupling between this deformation to the spin follows at the next half order. GW170817 placed constraints on each of these PN orders, as detailed in Abbott et al. 8.
With the above in mind, the notion of dephasing can be made precise. In the stationary phase approximation, the frequency-space GW phase, $\Psi$, can be written as [e.g. 142, 102]
@@ IV.1.2 Tides: equilibrium
The equilibrium tide simply corresponds to the “$\Phi(t)=0$” portion of the dynamics. In this case, the quasiharmonic time dependence of (13) drops out, and our interest shifts to bulk, geometric deformations of the stellar surface (“zero-frequency oscillations”). The extent to which the stellar interior is suspectible to a time-independent tidal potential can be encapsulated by the (shape) Love numbers; effectively, much like the $Q_{\alpha}$ defined previously, these measure the extent of orthogonality between the stellar fluid (or solid) and some angular portion of the tidal field (i.e. $\ell,m^{\prime}$). In general, therefore, there is an infinite set of Love numbers, though often in the literature one will find the term “Love number” just to mean the quadrupolar, $m^{\prime}=0$ Love number. In the context of compact binaries in GR, Damour 104 first quantified how tidal Love numbers influence inspiral. Note also that if the star is static and axisymmetric at background order, the index $m^{\prime}$ falls out of the equations and thus the static, or even effective, Love numbers are characterised by a single harmonic number $\ell$. These reduced Love numbers are typically denoted $k_{\ell}$.
@@
which is the same $U$ from (13). Note that the multipole moments $I_{\ell m}$ are also implicitly defined at some PN order; see Section V in Thorne 401. At Newtonian order, $I_{\ell m}$ is defined through an integral over the mass-density weighted by a spherical harmonic and radius to power $\ell$:
@@
where the (axisymmetric) mass ($M_{\ell}$) and current ($S_{\ell})$ multipole moments can be defined either via the Thorne 401 or Geroch-Hansen [178] formulae. These formulations are totally equivalent by a theorem due to Gürsel 174. In the above we have the odd and even parity sectors of contractions of the trace-free Weyl tensor [e.g. 108, 297]. Using the schemes introduced by Pappas and Sotiriou 299, Suvorov and Melatos 388 to generalise the Geroch-Hansen definitions to theories beyond GR, one could try to extend the correspondence (38) to some other theory of gravity. This has not yet been attempted.
Figure 7: The main radial gravitational potential $A(R)$ — the lapse function of the effective-one-body spacetime over which the dynamics of a test particle describes the dynamics of the real binary system [71, 72] — as a function of the radial coordinate $R=a$ with $M=M_{A}+M_{B}$ (for this Figure only) in a tidal-EOB model. The binary black hole sector of this model (solid black) shows a deviation from the lapse of the Schwarzschild spacetime (dotted), indicating the dynamics of a test particle on the effective spacetime is different from that of a (plunging) orbit around a Schwarzschild black hole due to the finite symmetric mass ratio $\nu=M_{A}M_{B}/M^{2}$ of the binary (which should not be confused with the our use of $\nu$ as stellar spin used elsewhere). For the neutron star case, the gravitational potential including up to next-to-next-to-leading-order tidal effects (i.e. up to 7 PN; green dashed) is shown together with the result resumming the gravitational-self-force information up to 7.5 PN [61, red]. The specific equal-mass neutron-star binary considered here has a tidal deformability (equation 39) of $\kappa_{2}^{T}\approx 73.55$. From Bernuzzi et al. 55 with permission.
Tides effectively enhance the “attraction” between the components of the binary, leading to earlier merger relative to binary black-holes or point-particles. The tidal dynamics are elegantly described by a radial interaction potential in the effective-one-body (EOB) formalism [62, 55, 187]. PN models of tides in the EOB formalism were initially verified against numerical simulations by Baiotti et al. 43, Bernuzzi et al. 56, who demonstrated that the models can be improved with such calibrations but that it is not possible to accurately describe tides close to merger with a PN model. Such a restriction was lifted by Bernuzzi et al. 55 through resumming techniques. A schematic depiction of this attractive feature is shown in Figure 7, where the red curve describes the radial potential for a neutron-star binary with tidal resumming. A significant drop in the potential occurs at larger radius, indicating the stronger attraction felt by the binary than the equivalent binary black hole. The dominant term of the tidal imprints on waveform is delivered by the quadrupole quantity [107]
@@
for $\Lambda_{A}=\frac{2}{3}k_{2}^{A}(R_{A}/M_{A})^{5}$ with the same definition for object $B$ [see also 185, 143, 186, 106]. Note $k_{2}$ is defined in expression (34), effectively being the constant of proportionality between the quadrupole moment and the quadrupolar tidal potential (appropriately generalised to GR via the Weyl tensor; though cf. Footnote 4).
@@ IV.1.3 Tides: dynamical
Pulsation modes inside compact stars, tidally-forced or otherwise, are generally characterised according to the nature of the restoring force that ultimately damps the oscillations. The theory of mode oscillations is an active and rich area of research, which we cannot do justice in this Review. We therefore refer the reader to, e.g. Andersson and Kokkotas 31, Kokkotas et al. 216, Pratten et al. 319, Andersson 24, though a few QNM groups are particularly important in the context of GWs and tides, so we introduce them briefly.
@@
Modes that are instead restored by buoyancy are referred to as $g$-modes (i.e. “gravity” modes, not to be confused with GWs). The $g$-modes, in contrast to $f$-modes, do depend sensitively on the internal composition of the star (see Sec. III.3) and thus may be able to reveal a different kind of information. Depending on the chemical composition of the star, the $g$-modes tend to follow different relations as described by Kuan et al. 228. That is to say, whether the EOS is purely hadronic or hybrid (for instance) has a strong impact on the scaling of the modes with the mean density, temperature, and other, microphysical parameters. Generally speaking, the $g$-mode spectrum will be influenced by both composition and temperature (entropy) gradients. The former can be calculated from a EOS that provides the speed of sound and $d\epsilon/dp$ in a tabulated form so that the Brunt-Väisälä frequency can be determined [11, 409].
@@
The $p$-modes, and $f$-mode in particular, are especially important for astrophysical processes involving neutron stars owing to their compactness, including the tides: it can be shown that the $f$-mode couples most strongly to the tidal potential $\nabla U$ out of all other modes (at least for astrophysical stars, e.g. without near break-up rotation rates or virial-strength magnetic fields). The $f$-modes tend to have a (linear) frequency in the range of $2\lesssim\omega_{f}/2\pi\lesssim 3$kHz depending on the EOS [see 218, for a review], and thus are primarily relevant in the late stages of an inspiral. The dephasing induced by the $f$-mode oscillation is a problem that has attracted considerable attention, where the effects are usually incorporated via the Love number as an effective dressing [187, 365, 253, 35, 366, 154]. Considering rapidly rotating stars leads to a more complicated picture, since the retrograde mode may actually come into resonance much earlier as the inertial-frame frequency of the mode is $\omega_{i}=\omega_{0}-m\Omega_{s}$ (equation 24; see Kuan and Kokkotas 226, 227).
@@
In Kuan et al. 230, the method presented in Kokkotas and Schafer 217 to simultaneously evolve the inspiral and mode amplitudes was extended to 3.5 PN order. In particular, the conservative dynamics, leading-order gravitational radiation, and mode excitations (see Tab. 2) are incorporated into a total system Hamiltonian. Solving the associated equations of motion allows one to quantify the above effects, including that of spin-modulations in mode excitation (deferred to Sec. IV.3). The aforementioned physics can also be incorporated into the effective Hamiltonian of the EOB formalism [187, 365], where spin can also be included [366]. We note that only modes whose pattern speed is along the orbital motion can be considerably excited while there is another set of modes rotating in the opposite direction. Hereafter, we always discuss modes belonging to the former class unless stated otherwise.
@@
Figure 8: Left panel: self-consistent mode amplitude evolutions of $f$- and $g_{1,2,3}$-modes due to tidal forcing, together with the dephasing (lower; see equation 31) associated with the $f$- (magenta) and $g$-modes (other colors; see plot legends) as functions of time relative to the merger, $t-t_{\rm mrg}$. The GW frequency ($f_{\rm gw}$) is overplotted as the green line, corresponding to the vertical axis on the right-hand axis. Right: Same information as per the left panel though instead as functions of $f_{\rm gw}$, with $t-t_{\rm mrg}$ overlaid as the green line, corresponding to the vertical, right-hand axis. The evolution is carried for a $1.35+1.35\,M_{\odot}$ binary neutron-star (i.e. $q=1$) with the generalised piecewise-polytropic approximation of the APR EOS from O’Boyle et al. 288 (the original version of the piecewise-polytropic approximation in Read et al. 331 has discontinuous $\tilde{\Gamma}$, preventing an accurate determination of $g$-modes’ tidal properties when $\delta$ is small) with a constant stratification $\delta=0.005$. The vertical lines mark the frequency of $g_{1,2,3}$-modes, which are 90, 60 and 48 Hz, respectively.
@@
On the other hand, if the neutron star has a spin, the frequencies of retrograde modes will be reduced [99, 243, 440, 153], pushing the onsets of resonance earlier in the inspiral. Of the most relevant effects in terms of waveform is the $f$-mode excitation in a neutron star possessing an anti-aligned spin with the orbit [190, 253, 366], as also demonstrated in numerical simulations [e.g. 128, 154]; effects of spin are described in more detail in Sec. IV.3.
Current analytic waveform models include either implicitly or explicitly dynamical tides to introduce some effective dressing. In particular, Andersson and Pnigouras 34 show that the Love number can be expressed as a sum of the contributions of various QNMs with the $f$-mode dominating. As the orbital frequency increases, the tidal interaction enters a dynamical regime, where the Love number varies with tidal frequency and is often referred to as an effective Love number [187, 365]. Another stream to model the dynamical tidal effects is to introduce amplification factors to the leading-order tidal effects to blend in higher-order PN contributions [106, 62, 283, 13]. Phenomenological fittings to numerical simulations have also been developed [120, 123, 122, 4] [see also 429].